# Symplectic Quantization: numerical results for the Feynman propagator on a 1+1 lattice and the theoretical relation with Quantum Field Theory

### v2

@@ VIII Conclusions and Perspectives
In this work we have presented the first numerical test of symplectic quantization, a new functional approach to quantum field theory [8, 9] which allows for an importance sampling procedure directly in Minkowski space-time. The whole idea, which parallels the one of stochastic quantization, is based on the assumption that fields have a dependence of an additional time parameter, the intrinsic time $\tau$, with respect to which conjugated momenta $\pi(x)$ are defined. Quantum fluctuations of the fields are sampled by means of a deterministic dynamics flowing along the new time $\tau$, which controls the internal dynamics of the system and is distinguished from the coordinate time of observers and clocks. Such a dynamics is generated by a generalized Hamiltonian where the original relativistic action plays the role of a potential energy part and therefore fluctuates naturally along the flow of $\tau$. This whole construction does not need any sort of rotation from real to imaginary time to be consistent and to efficiently allow the numerical sampling of field fluctuations. Furthermore, under the hypothesis of ergodicity, symplectic quantization allows to define a generalized microcanonical ensemble which represents a probabilistically well defined functional approach to quantum field theory. In Sec. VII we have shown that the microcanonical partition function corresponding to the symplectic quantization dynamics is simply connected by means of an integral transformation to the Feynman path integral, thus implying that also all the disconnected correlation functions measured from the symplectic quantization approach are connected by means of an integral transformation to quantum field theoretic correlations. This said, there is also another possible way to connect the microcanonical functionals studied in this work to the standard Feynman path integral. We have shown that it is possible to explicitly compute $\Omega[\hbar/z,J]$ in the continuum limit, where it turns out to be equivalent to a Minkowskian statistical mechanics theory with canonical weight $P[\phi]\propto\exp(zS[\phi]/\hbar)$. First of all, consistently with the results of our simulations, it must be noticed that this canonical probability is well defined for a Minkwoskian theory, provided that the potential is bounded from below, it is therefore a very promising tool to study non-perturbative problems in causal space-time. Second, the above canonical expression suggests that a punctual correspondence with standard quantum field theory can be drawn by analytically continuing along a suitable integration path the above weight in the complex $z$ plane [20, 21]. This last path looks very promising and is currently under investigation.
@@ Appendix A Numerical Algorithm
All numerical cacalculations in this paper have been performed using a siplitting algorithm of second order, which takes advance of the generalized Hamiltonian separability. Using the notation of Sec. II, the algorithm can be characterized as a map
$\Psi_{\delta\tau}:\phi(x,\tau),\pi(x,\tau)\leavevmode\nobreak\ \leavevmode\nobreak\ \longrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \phi(x,\tau+\delta\tau),\pi(x,\tau+\delta\tau),$ (83)
@@
Figure 7: Energy fluctuations $\delta E(\delta\tau)$ as a function of the timestep $\delta\tau$ of the numerical algorithm in the case of Minkowski metric and periodic boundary conditions. Energy conservation at the algorithmic precision, i.e. $\delta E(\delta\tau)\sim\delta\tau^{2}$, is fulfilled.
Figure 8: Energy fluctuations $\delta E(\delta\tau)$ as a function of the timestep $\delta\tau$ of the numerical algorithm in the case of Minkowski metric and fringe boundary conditions. Energy conservation at the algorithmic precision, i.e. $\delta E(\delta\tau)\sim\delta\tau^{2}$, is fulfilled.
$\pi(x,\tau+\delta\tau/2)$ $=\pi(x,\tau)+\frac{\delta\tau}{2}\cdot F[\phi(x,\tau)]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ x$
$\phi(x,\tau+\delta\tau)$ $=\phi(x,\tau)+\delta\tau\cdot\pi(x,\tau+\delta\tau/2)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ x$
$\pi(x,\tau+\delta\tau)$ $=\pi(x,\tau+\delta\tau/2)+\frac{\delta\tau}{2}\cdot F[\phi(x,\tau+\delta\tau)]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ x$
@@
$\delta E(\delta\tau)=\langle|E(\tau)/E_{0}-1|\rangle\leavevmode\nobreak\ \propto\leavevmode\nobreak\ \delta\tau^{2}$ (86)
We have verified that the bound in Eq. (86) is fulfilled by the fluctuations of both the Hamiltonian $E(\tau)=\mathbb{H}[\phi(x,\tau),\pi(x,\tau)]$ in the case of Minkowski metric with periodic boundary conditions and the total Hamiltonian (system + boundary layers) $\mathbb{H}_{f}[\phi(x,\tau),\pi(x,\tau)]$ in the case of fringe boundary conditions (See Eq. (39) and the following discussion for the definition of $\mathbb{H}_{f}[\phi,\pi]$). In Fig.7 and Fig.8 is shown the behavior of $\delta E(\delta\tau)$ as a function of $\delta\tau$ respectively for the case of periodic and fringe boundary conditions.

### v3

@@ 8 Conclusions and Perspectives
In this work we have presented the first numerical test of symplectic quantization, a new functional approach to quantum field theory GL21 ; G21 which allows for an importance sampling procedure directly in Minkowski space-time. The whole idea, which parallels the one of stochastic quantization, is based on the assumption that fields have a dependence of an additional time parameter, the intrinsic time $\tau$, with respect to which conjugated momenta $\pi(x)$ are defined. Quantum fluctuations of the fields are sampled by means of a deterministic dynamics flowing along the new time $\tau$, which controls the internal dynamics of the system and is distinguished from the coordinate time of observers and clocks. Such a dynamics is generated by a generalized Hamiltonian where the original relativistic action plays the role of a potential energy part and therefore fluctuates naturally along the flow of $\tau$. This whole construction does not need any sort of rotation from real to imaginary time to be consistent and to efficiently allow the numerical sampling of field fluctuations. Furthermore, under the hypothesis of ergodicity, symplectic quantization allows to define a generalized microcanonical ensemble which represents a probabilistically well defined functional approach to quantum field theory. In Sec. 7 we have shown that the microcanonical partition function corresponding to the symplectic quantization dynamics is simply connected by means of an integral transformation to the Feynman path integral, thus implying that also all the disconnected correlation functions measured from the symplectic quantization approach are connected by means of an integral transformation to quantum field theoretic correlations. This said, there is also another possible way to connect the microcanonical functionals studied in this work to the standard Feynman path integral. We have shown that it is possible to explicitly compute $\Omega[\hbar/z,J]$ in the continuum limit, where it turns out to be equivalent to a Minkowskian statistical mechanics theory with canonical weight $P[\phi]\propto\exp(zS[\phi]/\hbar)$. First of all, consistently with the results of our simulations, it must be noticed that this canonical probability is well defined for a Minkwoskian theory, provided that the potential is bounded from below, it is therefore a very promising tool to study non-perturbative problems in causal space-time. Second, the above canonical expression suggests that a punctual correspondence with standard quantum field theory can be drawn by analytically continuing along a suitable integration path the above weight in the complex $z$ plane Witten:2010cx ; witten2010newlookpathintegral . This last path looks very promising and is currently under investigation.
@@ Appendix A Numerical Algorithm
All numerical cacalculations in this paper have been performed using a siplitting algorithm of second order, which takes advance of the generalized Hamiltonian separability. Using the notation of Sec. 2, the algorithm can be characterized as a map
$\Psi_{\delta\tau}:\phi(x,\tau),\pi(x,\tau)~~\longrightarrow~~\phi(x,\tau+\delta\tau),\pi(x,\tau+\delta\tau),$ (83)
@@
(a) Energy fluctuations $\delta E(\delta\tau)$ as a function of the timestep $\delta\tau$ of the numerical algorithm in the case of Minkowski metric and periodic boundary conditions. Energy conservation at the algorithmic precision, i.e., $\delta E(\delta\tau)\sim\delta\tau^{2}$, is fulfilled.
(b) Energy fluctuations $\delta E(\delta\tau)$ as a function of the timestep $\delta\tau$ of the numerical algorithm in the case of Minkowski metric and fringe boundary conditions. Energy conservation at the algorithmic precision, i.e., $\delta E(\delta\tau)\sim\delta\tau^{2}$, is fulfilled.
Figure 6: Energy fluctuations $\delta E(\delta\tau)$ for different boundary conditions in the Minkowski metric.
$\pi(x,\tau+\delta\tau/2)$ $=\pi(x,\tau)+\frac{\delta\tau}{2}\cdot F[\phi(x,\tau)]~~~~~~~~\forall~x$
$\phi(x,\tau+\delta\tau)$ $=\phi(x,\tau)+\delta\tau\cdot\pi(x,\tau+\delta\tau/2)~~~\forall~x$
$\pi(x,\tau+\delta\tau)$ $=\pi(x,\tau+\delta\tau/2)+\frac{\delta\tau}{2}\cdot F[\phi(x,\tau+\delta\tau)]~~~\forall~x$
@@
$\delta E(\delta\tau)=\langle|E(\tau)/E_{0}-1|\rangle~\propto~\delta\tau^{2}$ (86)
We have verified that the bound in Eq. (86) is fulfilled by the fluctuations of both the Hamiltonian $E(\tau)=\mathbb{H}[\phi(x,\tau),\pi(x,\tau)]$ in the case of Minkowski metric with periodic boundary conditions and the total Hamiltonian (system + boundary layers) $\mathbb{H}_{f}[\phi(x,\tau),\pi(x,\tau)]$ in the case of fringe boundary conditions (See Eq. (39) and the following discussion for the definition of $\mathbb{H}_{f}[\phi,\pi]$). In Fig.6(a) and Fig.6(b) is shown the behavior of $\delta E(\delta\tau)$ as a function of $\delta\tau$ respectively for the case of periodic and fringe boundary conditions.
Appendix B Fringe boundaries
In this section we present results of a preliminary investigation into the sensitivity of the measured mass to the parameters governing the fringe boundary conditions. These numerical tests were conducted with the same parameters used for Fig. 5 in the main text, namely $m=1$, $\lambda=0.001$, $a=1$ and $L=128$, unless otherwise specified. Our findings are summarized in Table 1 and discussed below.
In the first place, we investigated the behavior of the measured mass upon varying the damping parameter $\varepsilon$, while keeping it constant across space. The top section of Table 1 shows that even by varying $\varepsilon$ over ten orders of magnitude (from $10^{-5}$ down to $10^{-15}$) we do not find evidence of any trend in the change of the measured mass, which therefore seems not affected by variations of $\varepsilon$. We also found that simply increasing the fringe region thickness $L_{\text{fringe}}$ from $L$ to $2L$ also makes apparently no difference.
We then investigated whether a spatially varying profile $\varepsilon(x)$ may influence the finite-size effects on the measured mass. We considered both a linear and an exponential decay for $\varepsilon(x)$ across the fringe layer. As shown in the middle section of Table 1, a linear decay profile seems to have no effect on the measured mass. On the contrary, by exploiting an exponential decay of the form $\varepsilon(x)=\exp(-cx)$ we find an encouraging signal: the measured mass value seems to have a trend towards the expected value as long as the decay becomes sharper (i.e., for larger $c$).
Finally, we made a very preliminary investigation on how the measured mass value depends on the ratio $L_{\text{fringe}}/L$, upon increasing $L$ at fixed $L_{\text{fringe}}$. From the bottom section of Table 1 we have some preliminary indication that decreasing the ratio $L_{\text{fringe}}/L$ there is a trend towards the expected value of the mass, but a more systematic investigation is clearly in order.
Table 1: Investigation of the dependence of the measured mass on fringe boundary condition parameters. The baseline from the paper is $L=128$, $L_{\text{fringe}}=L$, constant $\varepsilon=10^{-10}$, yielding a mass of $2.06\pm 0.05$.
Test Type Parameters Measured Mass
Constant $\varepsilon$ $\varepsilon=10^{-3}$ $2.09\pm 0.04$
($L=128,L_{\text{fringe}}=L$) $\varepsilon=10^{-5}$ $2.06\pm 0.03$
$\varepsilon=10^{-10}$ $2.06\pm 0.05$
$\varepsilon=10^{-15}$ $1.97\pm 0.04$
Varying $\varepsilon(x)$ form Linear Decay $2.31\pm 0.06$
($L=128,L_{\text{fringe}}=L$) Exponential Decay, $c=0.05$ $2.22\pm 0.07$
Exponential Decay, $c=0.5$ $2.13\pm 0.06$
Exponential Decay, $c=1.0$ $1.99\pm 0.04$
Varying Lattice Size $L=128,L_{\text{fringe}}=L,\varepsilon=10^{-10}$ $2.06\pm 0.05$
($L_{\text{fringe}}=128$) $L=256,L_{\text{fringe}}=L/2,\varepsilon=10^{-10}$ $1.91\pm 0.02$
$L=128,L_{\text{fringe}}=L,\varepsilon=10^{-15}$ $1.97\pm 0.04$
$L=256,L_{\text{fringe}}=L/2,\varepsilon=10^{-15}$ $1.91\pm 0.02$